# Always Keep Your Promises: A Model-Agnostic Attribution Algorithm for Neural Networks

### v3

@@ Abstract
Inability to precisely understand neural network outputs is one of the most severe issues limiting the use of AI in multiple domains, from science and medicine to high-stakes decision and regulatory models. Layer-wise Relevance Propagation (LRP) is an established explainability method that addresses some of these limitations, but widespread adoption has not been possible because existing implementations must be coupled with individual model architectures, rendering them impractical for a rapidly evolving model space. We present DynamicLRP, a lightweight and flexible method for performing LRP on any neural network with provably minimal overhead. To achieve this, we introduce a novel graph search mechanism called the “Promise System”, which repurposes deep learning computation graphs for non-gradient computations and is implemented at the level of primitive tensor operations using standard automatic differentiation libraries. We demonstrate that DynamicLRP matches or surpasses specialised implementations in attribution quality across vision and language tasks and remains efficient for models at the billion-parameter scale. Notably, our implementation achieved 99.99% operation coverage across 31,465 computation nodes from 15 diverse architectures, without any architecture-specific modifications. This is the first truly model-agnostic LRP solution, enabling high-quality neural network attribution across the full spectrum of modern AI architectures. All code is available at https://github.com/keeinlev/dynamicLRP.
@@ 1 Introduction
Modern neural networks, in particular massive transformer-based foundation models Gemma Team (2025); Dosovitskiy et al. (2021); Grattafiori et al. (2024), offer unprecedented ability to map complex, unstructured inputs to accurate outputs. Inability to understand this mapping severely limits use of these powerful models in all domains where the reverse mapping is wanted. Pertinent domains include, but are not limited to: natural science, social science, medicine, law, and finance.
While a variety of methods for Explainable AI (XAI) do exist, each has substantive drawbacks. The most well-known methods include perturbation-based approaches such as SHAP Lundberg and Lee (2017) and LIME Ribeiro et al. (2016). These methods are thoroughly unsuitable for use in massive models, with SHAP explanations becoming more misleading with input cardinality Huang and Marques-Silva (2024), and LIME suffering from an arbitrary definition of sensitivity to locality. These methods are further troubled by the exponential computation incurred proportional to model complexity Covert et al. (2020). Gradient-based methods Simonyan et al. (2014); Sundararajan et al. (2017); Smilkov et al. (2017) are fragile, suffering from saturation, shattering, and assumptions of linearity Ancona et al. (2018); Samek et al. (2021). Class Activation Mapping Zhou et al. (2016); Selvaraju et al. (2017) lacks granular precision. In contrast, Layer-wise Relevance Propagation (LRP), a backpropagation-based method, Bach et al. (2015) provides a principled, efficient alternative by decomposing predictions into input contributions via strict conservation properties Montavon et al. (2017), avoiding the pitfalls of local sensitivity Samek et al. (2021) and proving effective in diverse domains Iwana et al. (2019); Böhle et al. (2019); Sun et al. (2021).
Despite its theoretical advantages, LRP remains constrained by practical limitations tied to module-specific rules, severely limiting adaptability to rapidly evolving architectures like Transformers Achtibat et al. (2024); Otsuki et al. (2024). Modern libraries like iNNvestigate Alber et al. (2018), Captum Kokhlikyan et al. (2020), and notably Zennit Anders et al. (2021) attempt to mitigate this with automated differentiation, but remain module-centric: new architectures typically need new composite rules and custom configurations. Even specialized advances like AttnLRP Achtibat et al. (2024) (built on Zennit) require manual adaptation for new model updates (e.g., Llama 2 vs. 3), creating friction and overhead for continued use.
@@
To address these limitations, we introduce DynamicLRP, which shifts the focus from a module-centric to a primitive-based formulation by redefining the propagation rules at the level of atomic tensor operations (e.g., addition, matrix multiplication). Analogous to generalized gradient computation, DynamicLRP implements LRP rules for 47 primitives and propagates relevance directly through the computation graph, making the method inherently model-agnostic and free of model-specific configuration (Figure 1). This design is enabled by the Promise System, a novel deferred propagation mechanism that resolves the challenge of missing operand information during non-gradient computations within auto-differentiation frameworks (Sec 4.2). Lastly, we show that our approach matches or exceeds specialized implementations in attribution quality (Sec 5) across popular portable vision (ViT-b-16 Dosovitskiy et al. (2021), VGG16 Simonyan and Zisserman (2015)) and language (LLaMA Grattafiori et al. (2024), BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), T5-Flan Lee (2023)) models while maintaining efficiency on models into the billion-parameter scale.
2 Related Work
2.1 Explainability Methods
Explainable AI (XAI) algorithms fundamentally seek to decompose model predictions into input feature contributions, yet differ dramatically in their mathematical foundations and computational approaches. Gradient-based methods leverage the chain rule to compute $\frac{\partial f}{\partial x}$ as attribution signals Simonyan et al. (2014); Sundararajan et al. (2017); Smilkov et al. (2017), but suffer from gradient saturation, gradient shattering, and linearity assumptions that fail to capture complex nonlinear interactions Ancona et al. (2018); Samek et al. (2021). Perturbation-based approaches like SHAP Lundberg and Lee (2017) and LIME Ribeiro et al. (2016) directly measure prediction sensitivity through game-theoretic frameworks or local surrogates, yet incur exponential computational costs, making them intractable for high-dimensional data Covert et al. (2020). Class Activation Mapping methods Zhou et al. (2016); Selvaraju et al. (2017) leverage architectural constraints for spatial localization but provide coarse-grained explanations lacking principled mathematical foundations for complex architectures. This positions LRP as a strong explainability method when used in the correct domains, as it does not suffer from these pitfalls due to its tight rooting in the model architecture and efficient computation requirements of a single backward pass Samek et al. (2021).
2.2 Modern LRP Implementations
LRP has evolved significantly since its inception, with modern implementations striving to achieve greater flexibility and efficiency. Several libraries, including iNNvestigate Alber et al. (2018), Captum Kokhlikyan et al. (2020), and Zennit Anders et al. (2021), provide implementations of LRP alongside other interpretability methods. Among these, Zennit stands out for its comprehensive support for rule-based methods in PyTorch, offering a modular framework that leverages the automatic differentiation machinery to backpropagate relevances intead of gradients. Another recent advancement in the field that is close to our work is AttnLRP Achtibat et al. (2024), which derived new propagation rules suitable for attention modules and provides an implementation named LxT (LRP eXplains Transformers) built on top of Zennit. Note that at the time of writing, the list of architectures currently supported by LxT is limited to Llama 2/3, Gemma 3, Qwen 2/3, BERT, GPT-2, and Vision Transformers. While LxT demonstrates the potential of extending LRP to modern architectures, it also illustrates the challenges of adapting module-centric designs and highlights the importance of a more flexible, architecture-agnostic approach.
While the implementation of new rules is claimed to be trivial in Zennit, there is nonetheless significant conceptual overhead in understanding how to correctly merge, split, and modify modules to fit within the framework, as well as additional management overhead due to the need for new configurations and codebase semantics when supporting other architectures. In particular, each model requires a custom, duplicated implementation to integrate the appropriate LRP propagation rules throughout the network. We build upon the foundation established by LxT and Zennit, acknowledging their flexibility while addressing the limitations of its module-centric approach. Specifically, adapting Zennit to previously unseen architectural patterns often requires the manual definition of new module mappings, a task that our operation-level approach seeks to automate.
@@ 3.1 Classic LRP via Taylor Approximation and LRP variants
The classic LRP formulation as presented by Bach et al. (2015) derives LRP using what is called the deep Taylor decomposition framework. This formulation is functional, treating the layers as mathematical operators, and thus it is devoid of the structural assumptions of any given architecture. By taking the first-order Taylor expansion of the function at some root point $x_{0}$ such that $f(x_{0})=0$, and invoking the proposed LRP relevance conservation property, they define a recursive expression for the relevance of neurons $x_{j}$ at layer $l$ w.r.t. neurons $x_{i}$ at layer l1l-1. For a detailed derivation, see Appendix A
The $\epsilon$-LRP Rule Assuming the bias contribution is negligible or handled separately (see Achtibat et al. (2024) for a detailed discussion on bias absorption), the input relevance can be computed by summing over all output neurons and introducing a stabilizing factor $\epsilon$ to properly handle vanishing denominators. For a linear layer $z_{j}=\sum_{i}x_{i}w_{ji}+b_{j}$ with $x_{0}=0$:
@@
The $\gamma$-LRP Rule Attributions can become noisy when using the $\epsilon$-LRP rule, particularly in scenarios where the denominator $z_{j}+\epsilon$ becomes dominated by the stabilizing factor $\epsilon$, resulting in locally insensitive relevance scores. The $\gamma$-LRP rule Montavon et al. (2019), can be used to address this by introducing the hyperparameter $\gamma$ to act as a multiplicative booster for positive neuron contributions:
@@
The Deep Taylor Decomposition framework has been recently used to derive propagation rules for all components in transformer-based architectures Vaswani et al. (2017). In their formulation, Achtibat et al. Achtibat et al. (2024) successfully derive rules for both bilinear matrix multiplications and softmax operations used in attention mechanisms. Our work builds on these insights to enable the reuse of these primitives for novel architectures.
@@ 4.1 Topological Propagation and the Problem of Missing Activations
Algorithm 1 Topological Traversal of Backward Graph
1:Input: Backward Graph $G=(V,E)$
2:Initialize $S$ as a stack containing all nodes with $indegree(v)=0$
3:Initialize $pending(v)=indegree(v)\quad\forall v\in V$
4:while $S$ is not empty do
5:  Pop node $v$ from $S$
7:  for $w\in outadj(v)$ do
8:    $pending(w)\leftarrow pending(w)-1$
9:    if $pending(w)==0$ then
10:     Push $w$ onto $S$
11:    end if
12:  end for
13:end while
@@ 4.2.3 Promise Deadlock
3:Initialize $\texttt{heuristic\_break\_queue}\leftarrow[\hskip 1.84995pt]$
4:Initialize $\texttt{input\_tracker}\leftarrow\{\texttt{node}:[\hskip 1.84995pt]\text{ for all nodes in graph}\}$
@@ 5 Experimental Setup
In this work, we aim to demonstrate that DynamicLRP produces faithful explanations for modern architectures of varying scale. In attempt to measure faithfulness, we adapt the evaluation strategies from Achtibat et al. (2024) using classification and exctractive question-answering tasks.
@@ 5.1 Classification Tasks
Input perturbation, as explored and advanced by Blücher et al. (2024), DeYoung et al. (2020), and Schulz et al. (2020), has become a widely-used standard for measuring attribution faithfulness. This method leverages the model itself as an oracle to validate the feature ranking generated by an attribution method. By systematically removing or replacing input features (e.g., tokens or pixels) and observing the degradation in performance with respect to a baseline, we can verify if a feature’s attributed “importance” is in-line with a model’s prediction. We quantify this using the Area Between Perturbation Curves (ABPC), ranking features by importance and iterating in both orders:
@@
We use the the base VGG16 model (Simonyan and Zisserman, 2015) evaluated on the ImageNette-320 dataset (Howard, 2019), and the base ViT-b-16 (Dosovitskiy et al., 2021) architecture on the CIFAR-10 dataset (Krizhevsky and Hinton, 2009). For perturbation, we occlude features by replacing them with a Gaussian-blurred version of the image ($\text{kernel size}=51$, $\sigma=20$). For ViT-b-16, we occlude the top-4 $16\times 16$ patches (pooled by max relevance) at each step. For VGG16, we occlude 1024 individual pixels per step.
@@
Following the experimental setup in Achtibat et al. (2024), we first use a finetuned Llama-3.2-1B model (Grattafiori et al., 2024) on the IMDb dataset (Maas et al., 2011) for sentiment analysis. We evaluate explanations by determining the relevance of input tokens to the predicted class. For language modelling, we use the base model on the English subset of the Wikipedia  [^1] Foundation dataset and frame the task as next-token prediction. We compute attribution for the ground-truth next token given a context window of 512 tokens. In both settings, we apply LRP-$\gamma$ with $\gamma_{linear}=1.0$ and compare against baseline attribution methods using MoRF and LeRF occlusion metrics after masking one token per step.
@@ 5.2 Extractive Question-Answering Tasks
Extractive Question-Answering (EQA) tasks differ structurally from classification: the model identifies a span within the input context as the answer. In this setting, the MoRF/LeRF strategy can be redundant, as the tokens constituting the predicted span are by definition necessary for the model to output that specific span. Therefore, we simplify the measurement to checking if the highest-ranked tokens are within the predicted answer span. We evaluate faithfulness via the accuracy of the top-1 attributed token falling within the label golden span and the model’s predicted span. We use finetuned versions of RoBERTa-large (Liu et al., 2019) and Flan-T5-large (Chung et al., 2022) on the SQuAD-v2 dataset (Rajpurkar et al., 2018; Chan et al., 2021; Lee, 2023). We include all answerable examples in the validation set, skipping only those where the model predicts an invalid span (start index > end index). We run LRP-$\gamma$ with $\gamma_{linear}=0.001$. We mask the relevance of tokens found in the question before taking the top-1 attributed token.
@@ 6.2 Operation Coverage
ViT-b-16 Dosovitskiy et al. (2021) Vision 779/779 16/16
Wav2Vec2-xls-r-300m Babu et al. (2021) Audio 2,021/2,022 18/19
Llama3.2-1B Grattafiori et al. (2024) Language 1,787/1,787 24/24
Mamba-130m Gu and Dao (2024) Language 3,421/3,421 26/26
@@ Appendix A LRP Derivation via Deep Taylor Decomposition
LRP redistributes the prediction score (relevance) backwards through a neural network. The challenge of faithful attribution lies in identifying a meaningful distribution rule for the relevance assigned to input $i$ by output neuron $j$, $R_{i\leftarrow j}$. We take advantage of the Deep Taylor Decomposition framework Montavon et al. (2017); Achtibat et al. (2024) to locally linearize and decompose operations into independent contributions, specifically by computing a first-order Taylor expansion of a function $f_{j}(x)$ at a reference point $x_{0}$:

### v4

@@ Abstract
Inability to precisely understand neural network outputs is one of the most severe issues limiting the use of AI in multiple domains, from science and medicine to high-stakes decision and regulatory models. Layer-wise Relevance Propagation (LRP) is an established explainability method that addresses some of these limitations, but widespread adoption has not been possible because existing implementations must be coupled with individual model architectures, rendering them impractical for a rapidly evolving model space. Our algorithm, DynamicLRP, is a lightweight and flexible method for performing LRP on any neural network with provably minimal overhead. To achieve this, we introduce a novel graph search mechanism called the “Promise System” which repurposes deep learning computation graphs for non-gradient computations and is implemented at the primitive tensor operation level using standard automatic differentiation libraries. We demonstrate that DynamicLRP matches or surpasses specialised implementations in attribution quality across vision and language tasks and remains efficient for models at the billion-parameter scale. Notably, our implementation achieved 99.99% operation coverage across 31,465 computation nodes from 15 diverse architectures, without any architecture-specific modifications. This is the first truly model-agnostic LRP solution, enabling high-quality neural network attribution across the full spectrum of modern AI architectures. All code is available at https://github.com/keeinlev/dynamicLRP.
@@ 1 Introduction
Modern neural networks, in particular massive transformer-based foundation models gemmateam2025gemma3; dosovitskiy2021vit; dubey2024llama3, offer unprecedented ability to map complex, unstructured inputs to accurate outputs. Inability to understand this mapping severely limits use of these powerful models in all domains where the reverse mapping is wanted. Pertinent domains include, but are not limited to: natural science, social science, medicine, law, and finance.
While a variety of methods for Explainable AI (XAI) do exist, each has substantive drawbacks. The most well-known methods include perturbation-based approaches such as SHAP lundberg2017unified and LIME ribeiro2016should. These methods are thoroughly unsuitable for use in massive models, with SHAP explanations becoming more misleading with input cardinality huang2024failings, and LIME suffering from an arbitrary definition of sensitivity to locality. These methods are further troubled by the exponential computation incurred proportional to model complexity covert2020understanding. Gradient-based methods simonyan2014deep; sundararajan2017axiomatic; smilkov2017smoothgrad are fragile, suffering from saturation, shattering, and assumptions of linearity ancona2018towards; samek2021xaireview. Class Activation Mapping zhou2016learning; selvaraju2017grad lacks granular precision. In contrast, Layer-wise Relevance Propagation (LRP), a backpropagation-based method, bach2015pixel provides a principled, efficient alternative by decomposing predictions into input contributions via strict conservation properties montavon2017explaining, avoiding the pitfalls of local sensitivity samek2021xaireview and proving effective in diverse domains iwana2019lrpexplainingcnns; bohle2019alzheimers; sun2021lrpfinetuning.
Despite its theoretical advantages, LRP remains constrained by practical limitations tied to module-specific rules, severely limiting adaptability to rapidly evolving architectures like Transformers achtibat2024attnlrp; otsuki2024lrpresidualconnections. Modern libraries like iNNvestigate alber2018innvestigate, Captum kokhlikyan2020captum, and notably Zennit anders2021zennit attempt to mitigate this with automated differentiation, but remain module-centric: new architectures typically need new composite rules and custom configurations. Even specialized advances like AttnLRP achtibat2024attnlrp (built on Zennit) require manual adaptation for new model updates (e.g., Llama 2 vs. 3), creating friction and overhead for continued use.
@@
To address these limitations, we introduce DynamicLRP, which shifts the focus from a module-centric to a primitive-based formulation by redefining the propagation rules at the level of atomic tensor operations (e.g., addition, matrix multiplication). Analogous to generalized gradient computation, DynamicLRP implements LRP rules for 47 primitives and propagates relevance directly through the computation graph, making the method inherently model-agnostic and free of model-specific configuration (Figure 1). This design is enabled by the Promise System, a novel deferred propagation mechanism that resolves the challenge of missing operand information during non-gradient computations within auto-differentiation frameworks (Sec 3.2). Lastly, we show that our approach matches or exceeds specialized implementations in attribution quality (Sec 4) across popular portable vision (ViT-b-16 dosovitskiy2021vit, VGG16 simonyan2015vgg) and language (LLaMA dubey2024llama3, BERT devlin2019bertpretrainingdeepbidirectional, RoBERTa liu2019roberta, T5-Flan lee2023flant5squadv2) models while maintaining efficiency on models into the billion-parameter scale.
@@ 2.1 Classic LRP via Taylor Approximation and LRP variants
The classic LRP formulation as presented by bach2015pixel derives LRP using what is called the deep Taylor decomposition framework. This formulation is functional, treating the layers as mathematical operators, and thus it is devoid of the structural assumptions of any given architecture. By taking the first-order Taylor expansion of the function at some root point $x_{0}$ such that $f(x_{0})=0$, and invoking the proposed LRP relevance conservation property, they define a recursive expression for the relevance of neurons $x_{j}$ at layer $l$ w.r.t. neurons $x_{i}$ at layer l1l-1. For a detailed derivation, see Appendix A
The $\epsilon$-LRP Rule Assuming the bias contribution is negligible or handled separately (see achtibat2024attnlrp for a detailed discussion on bias absorption), the input relevance can be computed by summing over all output neurons and introducing a stabilizing factor $\epsilon$ to properly handle vanishing denominators. For a linear layer $z_{j}=\sum_{i}x_{i}w_{ji}+b_{j}$ with $x_{0}=0$:
@@
The $\gamma$-LRP Rule Attributions can become noisy when using the $\epsilon$-LRP rule, particularly in scenarios where the denominator $z_{j}+\epsilon$ becomes dominated by the stabilizing factor $\epsilon$, resulting in locally insensitive relevance scores. The $\gamma$-LRP rule montavon2019lrp, can be used to address this by introducing the hyperparameter $\gamma$ to act as a multiplicative booster for positive neuron contributions:
@@
The Deep Taylor Decomposition framework has been recently used to derive propagation rules for all components in transformer-based architectures vaswani2017attention. In their formulation, Achtibat et al. achtibat2024attnlrp successfully derive rules for both bilinear matrix multiplications and softmax operations used in attention mechanisms. Our work builds on these insights to enable the reuse of these primitives for novel architectures.
@@ 3.2.3 Promise Deadlock
3:Initialize $\texttt{heuristic\_break\_queue}\leftarrow[\hskip 1.99997pt]$
4:Initialize $\texttt{input\_tracker}\leftarrow\{\texttt{node}:[\hskip 1.99997pt]\text{ for all nodes in graph}\}$
@@ 4 Experimental Setup
In this work, we aim to demonstrate that DynamicLRP produces faithful explanations for modern architectures of varying scale. In attempt to measure faithfulness, we adapt the evaluation strategies from achtibat2024attnlrp using classification and exctractive question-answering tasks.
@@ 4.1 Classification Tasks
Input perturbation, as explored and advanced by blucher2024pixelflipping, deyoung2020eraserbenchmarkevaluaterationalized, and schulz2020abpc, has become a widely-used standard for measuring attribution faithfulness. This method leverages the model itself as an oracle to validate the feature ranking generated by an attribution method. By systematically removing or replacing input features (e.g., tokens or pixels) and observing the degradation in performance with respect to a baseline, we can verify if a feature’s attributed “importance” is in-line with a model’s prediction. We quantify this using the Area Between Perturbation Curves (ABPC), ranking features by importance and iterating in both orders:
@@
We use the the base VGG16 model (simonyan2015vgg) evaluated on the ImageNette-320 dataset (howard2019imagenette), and the base ViT-b-16 (dosovitskiy2021vit) architecture on the CIFAR-10 dataset (krizhevsky2009learning). For perturbation, we occlude features by replacing them with a Gaussian-blurred version of the image ($\text{kernel size}=51$, $\sigma=20$). For ViT-b-16, we occlude the top-4 $16\times 16$ patches (pooled by max relevance) at each step. For VGG16, we occlude 1024 individual pixels per step.
@@
Following the experimental setup in achtibat2024attnlrp, we first use a finetuned Llama-3.2-1B model (dubey2024llama3) on the IMDb dataset (maas-IMDB) for sentiment analysis. We evaluate explanations by determining the relevance of input tokens to the predicted class. For language modelling, we use the base model on the English subset of the Wikipedia 111wikidump dataset and frame the task as next-token prediction. We compute attribution for the ground-truth next token given a context window of 512 tokens. In both settings, we apply LRP-$\gamma$ with $\gamma_{linear}=1.0$ and compare against baseline attribution methods using MoRF and LeRF occlusion metrics after masking one token per step.
@@ 4.2 Extractive Question-Answering Tasks
Extractive Question-Answering (EQA) tasks differ structurally from classification: the model identifies a span within the input context as the answer. In this setting, the MoRF/LeRF strategy can be redundant, as the tokens constituting the predicted span are by definition necessary for the model to output that specific span. Therefore, we simplify the measurement to checking if the highest-ranked tokens are within the predicted answer span. We evaluate faithfulness via the accuracy of the top-1 attributed token falling within the label golden span and the model’s predicted span. We use finetuned versions of RoBERTa-large (liu2019roberta) and Flan-T5-large (chung2022flant5) on the SQuAD-v2 dataset (rajpurkar2018squadv2; chan2021robertasquadv2; lee2023flant5squadv2). We include all answerable examples in the validation set, skipping only those where the model predicts an invalid span (start index > end index). We run LRP-$\gamma$ with $\gamma_{linear}=0.001$. We mask the relevance of tokens found in the question before taking the top-1 attributed token.
@@ 5.2 Operation Coverage
ViT-b-16 dosovitskiy2021vit Vision 779/779 16/16
Wav2Vec2-xls-r-300m babu2021xlsr Audio 2,021/2,022 18/19
Llama3.2-1B dubey2024llama3 Language 1,787/1,787 24/24
Mamba-130m gu2024mamba Language 3,421/3,421 26/26
@@ 7 Conclusion
Acknowledgements
The authors thank Julie Walsh, Dan Shonfeld, and Mirela Gondor from Scotiabank for their support and encouragement throughout this project. This work was made possible in part by a generous compute donation from Lambda Labs to PMA.
@@ Appendix A LRP Derivation via Deep Taylor Decomposition
LRP redistributes the prediction score (relevance) backwards through a neural network. The challenge of faithful attribution lies in identifying a meaningful distribution rule for the relevance assigned to input $i$ by output neuron $j$, $R_{i\leftarrow j}$. We take advantage of the Deep Taylor Decomposition framework montavon2017explaining; achtibat2024attnlrp to locally linearize and decompose operations into independent contributions, specifically by computing a first-order Taylor expansion of a function $f_{j}(x)$ at a reference point $x_{0}$:
@@ Appendix B Algorithms
Algorithm 2 Topological Traversal of Backward Graph
1:Input: Backward Graph $G=(V,E)$
2:Initialize $S$ as a stack containing all nodes with $indegree(v)=0$
3:Initialize $pending(v)=indegree(v)\quad\forall v\in V$
4:while $S$ is not empty do
5:  Pop node $v$ from $S$
7:  for $w\in outadj(v)$ do
8:   $pending(w)\leftarrow pending(w)-1$
9:   if $pending(w)==0$ then
10:     Push $w$ onto $S$
11:   end if
12:  end for
13:end while